3/4 divided by 8= x divided by 10and 2/3
Answer by stanbon(75887)

3/4 divided by 8= x divided by 10and 2/3
---------------------------
(3/4)/8 = x/(32/3)
------
x = (3/32)(32/3) = 1
============================
Cheers,
Stan H.
====================